factor
a^6n - B^6n

Answer by fractalier(2101)

a^6n - b^6n
This is the diference of squares...once we factor that, we then have the sum and difference of cubes and we can factor those too...here goes...
a^6n - b^6n =
(a^3n + b^3n)(a^3n - b^3n) =
(a^n + b^n)(a^2n - a^nb^n + b^2n)((a^n - b^n)(a^2n + a^nb^n + b^2n)